# Diffusion Models at the Drug Discovery Frontier: A Review on Generating Small Molecules versus Therapeutic Peptides

### v1

@@ Abstract
Diffusion models have emerged as a leading framework in generative modeling, showing significant potential to accelerate and transform the traditionally slow and costly process of drug discovery. This review provides a systematic comparison of their application in designing two principal therapeutic modalities: small molecules and therapeutic peptides. We analyze how a unified framework of iterative denoising is adapted to the distinct molecular representations, chemical spaces, and design objectives of each modality. For small molecules, these models excel at structure-based design, generating novel, pocket-fitting ligands with desired physicochemical properties, yet face the critical hurdle of ensuring chemical synthesizability. Conversely, for therapeutic peptides, the focus shifts to generating functional sequences and designing de novo structures, where the primary challenges are achieving biological stability against proteolysis, ensuring proper folding, and minimizing immunogenicity. Despite these distinct challenges, both domains face shared hurdles: the need for more accurate scoring functions, the scarcity of high-quality experimental data, and the crucial requirement for experimental validation. We conclude that the full potential of diffusion models will be unlocked by bridging these modality-specific gaps and integrating them into automated, closed-loop Design-Build-Test-Learn (DBTL) platforms, thereby shifting the paradigm from chemical exploration to the targeted creation of novel therapeutics.
@@ 1.1 The Bottleneck of Drug Discovery and the Rise of Generative AI
Traditional drug discovery pipelines, reliant on high-throughput screening and combinatorial chemistry, are characterized by prolonged development timelines, high attrition rates, and enormous costs. The entire process from target identification to market approval typically spans 10-15 years[1], with the clinical development phase alone requiring a median of 8.3 years[2]. Despite decades of optimization, clinical success rates remain discouragingly low, with only approximately 7.9% of drug candidates entering Phase I trials ultimately receiving regulatory approval[3], though these rates vary significantly across therapeutic areas and have shown dynamic fluctuations throughout the 21st century[4, 5, 6]. Recent advances in cell and gene therapies have demonstrated distinct success rate profiles, offering new prospects for durable treatments[7]. The financial burden is staggering: while historical estimates reached \$2.6 billion per approved drug[8], more recent analyses suggest mean development costs of approximately \$879 million based on 2000-2018 data[9], though costs continue to escalate with increasingly complex trial designs and regulatory requirements as evidenced by record-breaking FDA approval trends[10].
The vast chemical space, estimated to contain $10^{60}$ drug-like molecules[11, 12], remains largely unexplored through conventional screening approaches. This estimation, originally derived from molecules up to 30 heavy atoms constructed from organic elements[11], has been supported by systematic enumeration studies such as the GDB-17 database containing 166 billion molecules[13, 14] and recent explorations of peptide/peptoid chemical space[12]. More conservative estimates suggest approximately $10^{33}$ molecules strictly adhering to Lipinski’s rule of five[15], **yet even this reduced scope represents a vast and largely unsampled space.** Generative Artificial Intelligence (AI) offers a paradigm shift, moving from merely screening existing compounds to creating entirely novel molecules tailored to specific needs. Early generative models like Variational Autoencoders (VAEs)[16], Generative Adversarial Networks (GANs)[17], and Flow-based models[18, 19, 20] laid the groundwork but faced limitations in generation quality, training stability, and mode collapse issues. VAEs often produced blurry outputs due to reconstruction-latent loss trade-offs, while GANs suffered from training instability and mode collapse, and flow-based models encountered computational efficiency challenges[21, 22, 23].
@@ 1.2 The Emergence of Diffusion Models
Diffusion models have recently emerged as a highly successful framework in generative modeling, demonstrating competitive and robust capabilities in generating high-quality, diverse samples compared to previous approaches [24]. Their core idea involves a two-step process: a forward diffusion process that incrementally adds Gaussian noise to data according to a predefined variance schedule until it becomes pure noise, and a reverse denoising process where a trained neural network learns to iteratively denoise samples, effectively generating new data from random noise [25]. Mathematically, the forward process can be expressed as $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$, while the reverse process is learned as $p_{\theta}(x_{t-1}|x_{t})$ [25, 24].
The success of diffusion models spans multiple domains beyond molecular design. In computer vision, models like DALL-E 2 [26], Stable Diffusion [27], and Midjourney have dramatically advanced image generation, producing photorealistic images from text descriptions with remarkably high quality and controllability [28]. In audio synthesis, diffusion models have achieved state-of-the-art performance in speech generation [29, 30], music synthesis [31], and audio enhancement [32]. In video generation, recent models can create coherent, high-resolution videos [33]. In natural language processing, discrete diffusion models have been explored for controllable text generation [34], though autoregressive models remain dominant. The key to this cross-domain success lies in the flexibility of the diffusion framework: it can be adapted to different data modalities (continuous, discrete, structured) by designing appropriate noise processes and network architectures [24, 35]. The successful application of diffusion models in bioinformatics, covering areas from protein to small-molecule design, suggests their high potential in computational biology [36].
This remarkable versatility makes diffusion models particularly attractive for molecular design, where data exhibits diverse characteristics—continuous 3D coordinates, discrete atom types, graph structures, and sequential patterns. The success in other domains provides strong evidence that diffusion models can learn complex, high-dimensional distributions and generate high-quality, diverse samples [25, 24]. Moreover, techniques developed for image generation, such as classifier-free guidance [37] for controllable generation and latent diffusion [27] for computational efficiency, can be adapted to molecular design. Recent works have successfully applied these techniques to molecular discovery, including guided diffusion for inverse molecular design [38], evolutionary optimization for 3D multi-objective molecular generation [39], and de novo drug design [40], accelerating progress in the field. These models hold significant promise for advancing molecular sciences by creating diverse, valid, and novel molecular structures with desired properties.
@@ 1.3 Scope and Structure of This Review
Small molecules constitute a substantial portion of approved drugs. Recent FDA approval data from 2023-2024 indicate that small molecule drugs (new molecular entities, NMEs) accounted for approximately 55-69% of novel therapeutic approvals [41, 42, 43, 44]. In 2023, the FDA approved 55 new medications consisting of 17 biologics license applications and 38 NMEs, with small molecules representing approximately 55% of total approvals [41]. This approval trend continued in 2024, with 50 NMEs approved, further demonstrating the continued importance of small molecule drugs in modern therapeutics [42]. Small molecules typically have molecular weights below 900 Da, are orally bioavailable, can penetrate cells to target intracellular proteins, and are relatively cost-effective to manufacture. They have been successfully applied to a wide range of diseases, from infectious diseases (antibiotics, antivirals) to chronic conditions (cardiovascular drugs, diabetes medications) to cancer (kinase inhibitors, chemotherapeutics). However, small molecules face significant limitations in targeting certain ”undruggable” proteins, such as those lacking well-defined binding pockets or those involving protein-protein interactions with large, flat interfaces [45, 46, 47]. These challenging targets have spurred interest in alternative modalities and advanced drug design approaches.
@@ 2.1 Representing Molecules for Diffusion
The choice of molecular representation is fundamental to the design of the diffusion process, as it dictates both the mathematical formulation of the noise process and the architecture of the denoising network [53, 54]. For small molecules, representations primarily fall into two categories. One approach utilizes graph-based representations, where molecules are encoded as graphs with atoms as nodes and bonds as edges [55, 56, 57], allowing diffusion to operate on features like discrete atom types or continuous latent embeddings [54]. An alternative and increasingly prevalent approach employs 3D coordinate-based representations, treating molecules as point clouds of atomic positions in Euclidean space [58, 59, 60]. This latter representation is particularly suited for structure-based drug design, as it naturally captures spatial relationships critical for protein-ligand interactions and necessitates the use of E(3) equivariant neural networks to handle rotational and translational symmetries [61, 62, 63, 64, 65].
@@ 2.2 The Mathematics of Diffusion: Forward and Reverse Processes
The diffusion process consists of two Markov chains [25, 24]. The forward process gradually corrupts data $x_{0}$ by adding Gaussian noise over $T$ timesteps: $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$, where $\beta_{t}$ is a variance schedule. A key property is that we can sample $x_{t}$ directly from $x_{0}$: $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I)$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ [24]. The reverse process learns to denoise: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$. The model is trained to predict either the noise $\epsilon$ added at each step or the denoised data $x_{0}$, by minimizing a variational lower bound on the log-likelihood [25, 24]. For molecular generation, this framework is adapted to handle both continuous (coordinates) and discrete (atom/bond types, amino acid sequences) variables [35, 74, 75], often requiring specialized noise processes and network architectures.
@@ 2.4 Comparison with Other Generative Approaches
To appreciate the advantages of diffusion models, it is instructive to compare them with other generative paradigms previously applied to drug design. Variational Autoencoders (VAEs), such as the Junction Tree VAE [78], learn a continuous latent space but often struggle with posterior collapse and may generate chemically invalid structures when interpolating between latent points [79, 80, 81]. Similarly, Generative Adversarial Networks (GANs) like MolGAN [82] can produce diverse molecules but are notoriously difficult to train [83], frequently suffering from mode collapse and instability, especially when adapted for discrete graph structures [84, 85]. Flow-based models such as MoFlow [86] offer an alternative by learning an invertible mapping to a simple base distribution; however, the invertibility constraint can limit their architectural expressiveness when modeling complex molecular graphs [87, 88]. Autoregressive models like GraphAF [89], which build molecules sequentially, have shown strong performance but can be slow and suffer from error propagation, where an early mistake compromises the entire structure [90, 91, 40, 92, 93, 94].
@@ 3.1 Datasets and Benchmarks for Small Molecule Generation
The development and evaluation of diffusion models for small molecule design rely heavily on large-scale, high-quality datasets. The most widely used benchmark is CrossDocked2020 [104], a dataset containing approximately 22.5 million docked poses from over 100,000 protein-ligand complexes derived from the PDB (Protein Data Bank) through a systematic docking procedure [105]. Each complex includes the 3D coordinates of the protein binding pocket (typically defined as residues within 6-10 Å of the ligand) and the bound ligand, along with docking scores as a proxy for binding affinity. CrossDocked2020 has become the de facto standard for evaluating structure-based drug design models [106, 107, 53], enabling direct comparison across different approaches including diffusion-based methods [40] and other generative AI techniques [108]. However, it has several acknowledged limitations: the docking scores are computational estimates rather than experimental measurements, the dataset is biased toward certain protein families (kinases and proteases are over-represented), and the ligands are primarily known drugs or drug-like molecules, limiting chemical diversity. These limitations have motivated ongoing efforts to develop more diverse and experimentally validated benchmarks for the field.
@@
A critical limitation across all datasets is the scarcity of experimentally validated binding affinity data [115, 116, 117]. While databases like BindingDB [118] and ChEMBL [119] contain millions of bioactivity measurements, only a small fraction include high-resolution 3D structures of protein-ligand complexes, limiting the ability to train and validate structure-based generative models. This data scarcity motivates the development of transfer learning and semi-supervised approaches [120, 121, 122, 123, 124] that can leverage large unlabeled datasets.
@@ 3.2 Structure-Based Drug Design (SBDD)
DiffSBDD [107] introduces an SE(3)-equivariant graph neural network architecture that jointly diffuses over atomic coordinates and discrete atom types. By incorporating pocket information through a joint graph representation of the pocket and the growing molecule, DiffSBDD achieves superior performance in generating molecules with favorable predicted binding affinities. On the CrossDocked2020 dataset, DiffSBDD generates molecules with a median Vina score of -7.5 kcal/mol, outperforming previous autoregressive and VAE-based approaches. Importantly, the model demonstrates the ability to generate molecules that form key interactions (hydrogen bonds, hydrophobic contacts) with critical pocket residues, as validated through molecular dynamics simulations.
@@
Building upon these foundational approaches, recent work has explored dual diffusion frameworks and pharmacophore-oriented generation. Huang et al. [77] introduced a dual diffusion model that enables both de novo 3D molecule generation and lead optimization, providing a unified framework for structure-based drug discovery. More recently, pharmacophore-oriented approaches [126] have emerged to incorporate explicit pharmacophore constraints during the diffusion process, enabling more efficient feature-customized drug discovery by directly controlling key molecular properties and interaction patterns.

### v2

@@ Abstract
Diffusion models have emerged as a leading framework in generative modeling, poised to transform the traditionally slow and costly process of drug discovery. This review provides a systematic comparison of their application in designing two principal therapeutic modalities: small molecules and therapeutic peptides. We dissect how the unified framework of iterative denoising is adapted to the distinct molecular representations, chemical spaces, and design objectives of each modality. For small molecules, these models excel at structure-based design, generating novel, pocket-fitting ligands with desired physicochemical properties, yet face the critical hurdle of ensuring chemical synthesizability. Conversely, for therapeutic peptides, the focus shifts to generating functional sequences and designing de novo structures, where the primary challenges are achieving biological stability against proteolysis, ensuring proper folding, and minimizing immunogenicity. Despite these distinct challenges, both domains face shared hurdles: the scarcity of high-quality experimental data, the reliance on inaccurate scoring functions for validation, and the crucial need for experimental validation. We conclude that the full potential of diffusion models will be unlocked by bridging these modality-specific gaps and integrating them into automated, closed-loop Design-Build-Test-Learn (DBTL) platforms, thereby shifting the paradigm from mere chemical exploration to the on-demand engineering of novel therapeutics.
@@ 0.1 The Bottleneck of Drug Discovery and the Rise of Generative AI
Traditional drug discovery pipelines, reliant on high-throughput screening, which involves the automated testing of large numbers of compounds, and combinatorial chemistry, a method for rapidly creating vast libraries of molecules, are characterized by prolonged development timelines, high attrition rates, and enormous costs. The entire process from target identification to market approval typically spans 10–15 years [1], with the clinical development phase alone requiring a median of 8.3 years [2]. Despite decades of optimization, clinical success rates remain discouragingly low, with only approximately 7.9% of drug candidates entering Phase I trials ultimately receiving regulatory approval [3], though these rates vary significantly across therapeutic areas and have shown dynamic fluctuations throughout the 21st century [4, 5, 6]. Recent advances in cell and gene therapies have demonstrated distinct success rate profiles, offering new prospects for durable treatments [7]. The financial burden is staggering: while historical estimates reached \$2.6 billion per approved drug [8], more recent analyses suggest mean development costs of approximately \$879 million based on 2000–2018 data [9], though costs continue to escalate with increasingly complex trial designs and regulatory requirements as evidenced by record-breaking FDA approval trends [10].
The vast chemical space, estimated to contain $10^{60}$ drug-like molecules [11, 12], remains largely unexplored through conventional screening approaches. This estimation, originally derived from molecules up to 30 heavy atoms constructed from organic elements [11], has been supported by systematic enumeration studies such as the GDB-17 database containing 166 billion molecules [13, 14] and recent explorations of peptide/peptoid chemical space [12]. More conservative estimates suggest approximately $10^{33}$ molecules strictly adhering to Lipinski’s rule of five, a set of physicochemical guidelines used to predict a compound’s potential for oral bioavailability [15], yet even this reduced scope represents a vast and largely unsampled space. Generative Artificial Intelligence (AI) offers a paradigm shift, moving from merely screening existing compounds to creating entirely novel molecules tailored to specific needs. This promise is not merely theoretical; the broader field of generative AI has already begun to deliver tangible results, with dozens of AI-designed small molecules advancing into human clinical trials and demonstrating the potential to significantly shorten discovery timelines and improve success rates [16, 17, 18]. Early generative models like Variational Autoencoders (VAEs) [19], Generative Adversarial Networks (GANs) [20], and Flow-based models [21, 22, 23] laid the groundwork but faced limitations in generation quality, training stability, and mode collapse issues. VAEs often produced blurry outputs due to the trade-off between reconstruction and latent loss, while GANs were susceptible to training instability and mode collapse, challenges extensively reviewed in the literature focused on adversarial networks [24]. Flow-based models, in turn, encountered computational efficiency limitations. The distinct trade-offs in performance, stability, and computational cost across these generative families have been systematically compared in several surveys [25, 26].
@@ 0.2 The Emergence of Diffusion Models
Diffusion models have recently emerged as a highly successful framework in generative modeling, demonstrating competitive and robust capabilities in generating high-quality, diverse samples compared to previous approaches [27]. Their core idea involves a two-step process: a forward diffusion process that incrementally adds Gaussian noise to data according to a predefined variance schedule until it becomes pure noise, and a reverse denoising process where a trained neural network learns to iteratively denoise samples, effectively generating new data from random noise [28, 27].
The success of diffusion models extends far beyond a single domain. They have achieved revolutionary breakthroughs in fields like computer vision (e.g., DALL-E 2, Stable Diffusion), audio synthesis, and natural language processing [29, 30, 31, 32], proving an exceptional ability to learn and generate high-quality samples from complex, high-dimensional data. This cross-domain success underscores the framework’s inherent flexibility, which makes it particularly attractive for molecular design, where data is inherently multimodal—encompassing continuous 3D coordinates, discrete atom types, graph structures, and sequential patterns. Moreover, key techniques pioneered for image generation, such as classifier-free guidance [33] for precise control and latent diffusion [30] for computational efficiency, have been successfully adapted to molecular design challenges [34, 35]. This powerful combination of generative fidelity and adaptability provides a strong foundation for using diffusion models to create diverse, valid, and novel therapeutics with desired properties.
@@ 0.3 Scope and Structure of This Review
Small molecules constitute a substantial portion of approved drugs. Recent FDA approval data from 2023 to 2024 indicate that small molecule drugs (new molecular entities, NMEs) accounted for approximately 55–69% of novel therapeutic approvals [36, 37, 38, 39]. In 2023, the FDA approved 55 new medications consisting of 17 biologics license applications and 38 NMEs, with small molecules representing approximately 55% of total approvals [36]. This approval trend continued in 2024, with 50 NMEs approved, further demonstrating the continued importance of small molecule drugs in modern therapeutics [37]. Small molecules typically have molecular weights below 900 Da, are orally bioavailable, can penetrate cells to target intracellular proteins, and are relatively cost-effective to manufacture. They have been successfully applied to a wide range of diseases, from infectious diseases (antibiotics, antivirals) to chronic conditions (cardiovascular drugs, diabetes medications) to cancer (kinase inhibitors, chemotherapeutics). However, small molecules face significant limitations in targeting certain “undruggable” proteins—targets historically considered intractable for small-molecule intervention due to features like lacking well-defined binding pockets or those involving protein-protein interactions with large, flat interfaces [40, 41, 42]. These challenging targets have spurred interest in alternative modalities and advanced drug design approaches.
@@ 1.1 Representing Molecules for Diffusion
The choice of molecular representation is fundamental to the design of the diffusion process, as it dictates both the mathematical formulation of the noise process and the architecture of the denoising network [48, 49]. For small molecules, representations primarily fall into two categories. One approach utilizes graph-based representations, where molecules are encoded as graphs with atoms as nodes and bonds as edges [50, 51, 52], allowing diffusion to operate on features like discrete atom types or continuous latent embeddings [49]. An alternative and increasingly prevalent approach employs 3D coordinate-based representations, treating molecules as point clouds of atomic positions in Euclidean space [53, 54, 55]. This latter representation is particularly suited for structure-based drug design, as it naturally captures spatial relationships critical for protein-ligand interactions and necessitates the use of E(3) equivariant neural networks, architectures designed to respect the natural rotational and translational symmetries of 3D molecules, to handle rotational and translational symmetries [56, 57, 58, 59, 60].
@@ 1.2 The Mathematics of Diffusion: Forward and Reverse Processes
The diffusion process consists of two Markov chains [28, 27]. The forward process gradually corrupts data $x_{0}$ by adding Gaussian noise over $T$ timesteps: $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$, where $\beta_{t}$ is a variance schedule. A key property is that we can sample $x_{t}$ directly from $x_{0}$: $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I)$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ [27]. Here, $\beta_{t}$ is a small constant from a predefined variance schedule, controlling the amount of noise added at each step. By defining $\alpha_{t}=1-\beta_{t}$, the term $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ becomes a cumulative product that governs how much of the original signal $x_{0}$ is preserved at timestep $t$. As $t$ increases, $\bar{\alpha}_{t}$ decreases towards zero, signifying that the signal progressively fades into noise. Thus, $\bar{\alpha}_{t}$ can be intuitively understood as a measure of the signal-to-noise ratio at any given step. The reverse process learns to denoise: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$. The model is trained to predict either the noise $\epsilon$ added at each step or the denoised data $x_{0}$, by minimizing a variational lower bound on the log-likelihood [28, 27]. For molecular generation, this framework is adapted to handle both continuous (coordinates) and discrete (atom/bond types, amino acid sequences) variables [61, 70, 71], often requiring specialized noise processes and network architectures.
@@ 1.4 Comparison with Other Generative Approaches
To appreciate the advantages of diffusion models, it is instructive to compare them with the generative paradigms that were foundational and previously represented the state-of-the-art in drug design, such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), Flow-based models, and Autoregressive models. This comparison primarily focuses on the overarching generative framework rather than the underlying neural network architecture, as implementations of these paradigms often share powerful backbones like Graph Neural Networks (GNNs) or Transformers. The key distinctions, therefore, lie in their training objectives and generation mechanisms. For example, VAEs like the Junction Tree VAE [74] learn a continuous latent space but often struggle with posterior collapse and may generate chemically invalid structures [75, 76, 77]. Similarly, GANs like MolGAN [78] can produce diverse molecules but are notoriously difficult to train [79], frequently suffering from mode collapse and instability [80, 81]. Flow-based models such as MoFlow [82], constrained by their invertible architecture, can lack expressiveness for complex molecular graphs [83, 84]. Autoregressive models like GraphAF [85] can be slow and suffer from error propagation, where an early mistake compromises the entire structure [86, 87, 35, 88, 89, 90].
@@
These distinct generative mechanisms also lead to different trade-offs in computational cost, scalability, and interpretability. While frameworks like VAEs and GANs typically employ their backbone in a single-pass, feed-forward manner for generation, diffusion models operate iteratively, requiring hundreds or thousands of denoising steps. This iterative paradigm makes them more computationally intensive at inference time. Although the scalability of any single denoising step is governed by the underlying backbone (e.g., GNN or Transformer), the total generation cost is this value multiplied by the number of iterations. This computational overhead, however, is often offset by superior training stability, as diffusion models circumvent the notorious convergence issues and mode collapse that plague GANs. From an interpretability perspective, all these deep generative models face the ’black box’ challenge. Nevertheless, the step-by-step refinement process of diffusion models may offer unique, albeit still nascent, opportunities for mechanistic insight by allowing observation of the generative trajectory, though deciphering the rationale at each step remains an active area of research [27, 79, 35, 33].
@@ 2.1 Datasets and Benchmarks for Small Molecule Generation
The development and evaluation of diffusion models for small molecule design rely heavily on large-scale, high-quality datasets. The most widely used benchmark is CrossDocked2020 [100], a dataset containing approximately 22.5 million docked poses from over 100,000 protein-ligand complexes derived from the PDB (Protein Data Bank) through a systematic docking procedure [101]. Each complex includes the 3D coordinates of the protein binding pocket (typically defined as residues within 6–10 Å of the ligand) and the bound ligand, along with docking scores as a proxy for binding affinity. CrossDocked2020 has become the de facto standard for evaluating structure-based drug design models [102, 103, 48], enabling direct comparison across different approaches including diffusion-based methods [35] and other generative AI techniques [104]. However, it has several acknowledged limitations: the docking scores are computational estimates rather than experimental measurements, the dataset is biased toward certain protein families (kinases and proteases are over-represented), and the ligands are primarily known drugs or drug-like molecules, limiting chemical diversity. These limitations have motivated ongoing efforts to develop more diverse and experimentally validated benchmarks for the field.
@@
A critical limitation across all datasets is the scarcity of experimentally validated binding affinity data [111, 112, 113]. While databases like BindingDB [114] and ChEMBL [115] contain millions of bioactivity measurements, only a small fraction include high-resolution 3D structures of protein-ligand complexes. Beyond scarcity, the quality and heterogeneity of this data present a fundamental challenge to model reproducibility. Bioactivity measurements (e.g., Ki, Kd, IC50) are often aggregated from diverse assays with varying experimental conditions, introducing significant noise and inconsistencies. This lack of standardized data curation and reporting protocols directly undermines a model’s ability to learn robust structure-activity relationships, thereby compromising its generalization capabilities. Consequently, establishing rigorous data management standards is as critical as developing new algorithms for building truly predictive and reproducible generative models. This data challenge further motivates the development of transfer learning and semi-supervised approaches [116, 117, 118, 119, 120] that can leverage large unlabeled datasets while being robust to label noise.
@@ 2.2 Structure-Based Drug Design (SBDD)
DiffSBDD [103] introduces an SE(3)-equivariant graph neural network architecture that jointly diffuses over atomic coordinates and discrete atom types. By incorporating pocket information through a joint graph representation of the pocket and the growing molecule, DiffSBDD achieves superior performance in generating molecules with favorable predicted binding affinities. On the CrossDocked2020 dataset, DiffSBDD generates molecules with a median Vina score of -7.5 kcal/mol. This score serves as a computational estimate of binding affinity, where more negative values indicate a stronger predicted interaction, and this performance outperforms previous autoregressive and VAE-based approaches. Importantly, the model demonstrates the ability to generate molecules that form key interactions (hydrogen bonds, hydrophobic contacts) with critical pocket residues, as validated through molecular dynamics simulations.
@@
Building upon these foundational approaches, recent work has explored dual diffusion frameworks and pharmacophore-oriented generation. Huang et al. [73] introduced a dual diffusion model that enables both de novo 3D molecule generation and lead optimization, providing a unified framework for structure-based drug discovery. More recently, pharmacophore-oriented approaches [122] have emerged to incorporate explicit constraints on the pharmacophore—the essential three-dimensional arrangement of molecular features required for biological activity—during the diffusion process, enabling more efficient feature-customized drug discovery by directly controlling key molecular properties and interaction patterns.